Tech Staff Meeting Summary 7/23/02

Current Top Priorities are:Complete DBB, ANSS, start ESEE


2002 ANSS. Boards should be ready in a week or two. Will take 2 to 3 weeks for the Fiber "Fuse". Can begin replacing 2001 hardware without the "Fuse". Vaults should arrive by first week in August. Marked Tree is permitted. Mitch needs to put the station status list somewhere where people can see it. Does the new scream provide capability to get triggers? New firmware does not address problem of flash filling up when comm goes down and when in dual trigger/streaming mode. Gcf2ew module will be used for data streaming. Mitch will look into this. Reviewed proposed 2003 installations; approximately level effort though more distributed.

ESEE. Need contacts for blast sites ASAP since these will take the longest to set up and prep. Mitch needs to update the webpage.

GPS. Assume everything is running ok.

BB. GLAT still has problems. Repaired HALT and others. Now have a backlog of sensors to repair.

Old ANSS. We let these go until replacement with CMG5.

FDL. Instruments back and rehabbed. John needs to clear flash and install default config.

Oakridge. Greg will get with Munsey to schedule a scouting trip.

Dorman has a 2 day project for 3-4 MEQ/L28 to monitor a demolition in Greenville, MS. Will be done prior to 1 September. If it happens between about 14-26 August, Jim or Greg will need to do it, otherwise Dorman will.

Second Tier Priorities

Timing. Greg has an easy way to generate a datastream that goes high when lock is lost. Need a worm module to monitor. BB monitoring is more difficult. How do we automate the status stream or even get it back to Memphis for logging?

Still need to finish 4th channel up grade on about 1.5 nodes. Still need a degaining subroutine for ISIS.

Computer Room Power. Need to redo it when the generator is installed probably sometime in early September.

The original seisnets racks need new batteries.

DBB. Steve suggested a terminal server/watchdog to do dynamic resets when they hang. Has been working great though so this is backburner (need to order a bigger backburner and more pilot lights). ISTI worm module to monitor outages should best wait until RPC is running. Jim will check into getting the dish hung at NMAD and LNXT.


Greg was ruminating and came up with this todo list. Much of it that relates to other staff has been incorporated here. Til I figure out what to do with the rest of it, it is in unedited form.

MPHB and the box in the yard is up--sort of. Redirector software has been uninstalled. Mitch is working with testing several packages mentioned above under ANSS and they seem to be working. Should have MPHB online soon.

ET. Mitch shut down the lrnc heli-pages. John is focussing on new installs in east TN. The plan is to completely rebuild the NC stations when we can get time to get over and help. Then we'll put lrnc back online. Receiving all USNSN stations in Memphis. Steve replaced the computers for athn. Now there are no more bsd boxes. Worms running both in the EMA and on star mountain and streaming data. Still need to work out the details of getting data out of athn. Don't have the bandwidth for waveforms, but might send statrig messages and form trigger groups in Memphis. John found a bad preamp in in CPRT and is repairing. Repeater at FGTN is installed.

Monroe. Still want to. But when and where?

I-40 Bridge. Progressing nicely. Can probably do some installations later this year or early next year.

Need to do site visits of Lamont stations sometime this year.

3904/215E. Will expand into the closet and be used as a computer setup/testing lab.

Auburn. Wants to make some cores. They submitted a proposal. If funded will need our help getting a couple cores. Shouldn't be a big deal. A day to two maybe.

Calibrate S-13's.

SFTN ethernet redesign.

Battery charging station

renew 219 band licenses.

reattach 3904 gnd straps.

Repair fried SFTN k2 panel.

Miscellaneous notes

CVTN is running. Need one more visit to get comm going.

CBHT comm. Switch combined with dramatic k2ew config changes appeared to work. Steve said a new firmware upgrade may also solve the problem.

HDBT. Need a fence and gravel for ground cover.

NAIT. NWS does not know when the VPN will be installed. Mitch will ask the NWS Houston contact for more info.

CUET. Provided additional information to Pat Bohannan at TEMA regarding the software connections. He's working on the firewall issues (we hope).

SFTN. K2 and MSS100 are fine. No comm out the ethernet. Needs debugging.

TUMT. Mitch reconfigured all k2ew modules to be similar to CBHT (i.e. more robust with dropouts). Seems to have fixed TUMT. Now it seems to run out of open connections and only a reboot will fix (since you can't telnet in if there are no connections). Need to look into why sessions don't close.